OpenStudy (anonymous):

A barrel contains 109 gallons of water and is being drained at a constant rate of 7 gallons per hour. Write the equation that modles the number of gallons,g, after t hours.

OpenStudy (wolf1728):

After t hour, gallons in the barrel = 109 - (t *7) After 1 hour, gallons in the barrel = 109 - (1 *7) = 102 gallons

OpenStudy (wolf1728):

Or to make it just like the question: g = 109 - (t *7)